# Search for gravitationally lensed interstellar transmissions

### v1

@@ Abstract
We consider interstellar light transmission aided by a gravitational lens. We find that optimal reception efficiency occurs in lensing geometries where the transmitter, lens, and receiver are nearly aligned. We explore various signal detection strategies, employing both existing and emerging technologies. With this study, our understanding of interstellar power transmission via gravitational lensing has significantly progressed. We observe that detection signals from nearby stars, may leverage established photonics and optical engineering technologies, and networks of collaborative astronomical facilities. Our findings support the feasibility of interstellar power transmission via gravitational lensing, directly contributing to the ongoing optical SETI efforts.
@@ I Introduction
Interstellar power transmission is very challenging. Even for a collimated laser beam, the large distances involved result in a very small energy received. Consider, a diffraction-limited telescope with aperture $d_{\tt t}$ yielding the beam divergence of $\theta_{0}\simeq\lambda/d_{\tt t}=1.00\times 10^{-6}\,(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt t})$. When the signal reaches the receiver at distance $z$, the beam is expanded to a spot with the radius of $\rho_{*}\simeq z(\lambda/d_{\tt t})\simeq 3.09\times 10^{11}\,{\rm m}~{}(z/10\,{\rm pc})(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt t})$. As a result, in the case of a free-space laser power transmission in the vacuum, a telescope with the aperture $d_{\tt r}$ receives only a small fraction of the transmit power $P_{\tt T}$ given as $P_{\tt R}=P_{\tt T}{\pi({\textstyle\frac{1}{2}}d_{\tt r})^{2}}/{\pi\rho_{*}^{2}}$ or
@@
which yields photon flux of only $Q_{\tt R}=(\lambda/hc)P_{\tt R}\simeq 1.32\times 10^{-5}$  phot/s for $\lambda=1\,\mu$m. These signals are weak and require both 1) precise and stable transmitter pointing toward us to within $\rho_{\tt L}/z\simeq{\rm few}\,\mu$rad and 2) us having instruments capable of detecting such signals against the challenging optical backgrounds present in our stellar neighborhood.
Therefore, unless there is a significant effort to send us a message using a powerful, focused beam directly aimed at us, detecting faint and transient signals with optical SETI [^1] See details on the ongoing SETI efforts at https://en.wikipedia.org/wiki/Search_for_extraterrestrial_intelligence remains extremely challenging.
@@
This paper is organized as follows: In Section II, we present the wave-theoretical tools to describe light propagation in a gravity field. We discuss power transmission via a gravitational lens, introduce major noise sources, and evaluate detection sensitivity. In Section IV we discuss detection strategies. Our conclusions are presented in Section V.
@@ II EM waves in a gravitational field
We consider a nearby star with mass $M_{\tt L}$, radius $R_{\tt L}$, and its Schwarzschild radius of $r_{g}=2GM_{\tt L}/c^{2}$. We assume that an optical transmitter is placed in the focal region of this stellar gravitational lens at a distance $z_{0}>R^{2}_{\tt L}/2r_{g}=547.8\,(R_{\tt L}/R_{\odot})^{2}(M_{\odot}/M_{\tt L})~{}{\rm AU}$ from it.
@@
Figure 1: Beam propagation in a thin lens approximation showing transmitter, lens, annulus, and receiver plane.
We further assume that a CW laser optical transmitter is facing the lens and is capable of coherently illuminating an annular area around the lens with the radius of $b=\sqrt{2r_{g}z_{0}}>R_{\tt L}$ (see details in Turyshev (2024)). To do that, a diffraction-limited transmitter will have to form an annular beam of light with the mean diverging angle of $b/z_{0}=\sqrt{2r_{g}/z_{0}}\simeq 7.16\times 10^{-6}\,({b}/{R_{\tt L}})(R_{\tt L}/{R_{\odot}})({650\,{\rm AU}}/{z_{0}})~{}{\rm rad}$ and pointed toward the receiver behind the lens (see Fig. 1).
@@
When electromagnetic (EM) waves travel in the vicinity of the lens, its gravity causes the waves to scatter and diffract Turyshev (2017); Turyshev and Toth (2017). As a result, after passing by the lens and at a larger distance from it, $z\gg z_{0}$, the beam will be compressed and nearly collimated by the lens’ spherical aberration. If properly designed, the transmitted light will be deposited in the solar system. A solar system observer, looking back at the lensing star, will see the transmitted signal in the form of the Einstein ring. This focusing results in a major increase in the power density of the EM field deposited at the observer’s telescope (see discussion in Turyshev (2024)). Below we consider the physics of this effect.
II.1 Light amplification
@@
We introduce a lens-centric cylindrical coordinate system $(\rho,\phi,z)$ with its $z$-axis oriented along the unperturbed direction of the incident wave’s propagation (see Fig. 2), given by a unit vector $\boldsymbol{\mathrm{k}}$. We also introduce a light ray’s impact parameter, $\boldsymbol{\mathrm{b}}$, with $\boldsymbol{\mathrm{b}}\perp\boldsymbol{\mathrm{k}}$. In this coordinate system, the transmitter is positioned at $(0,-z_{0})$, while the receiver is on the plane with coordinates $(\boldsymbol{\mathrm{x}},z)$ and
${\boldsymbol{\mathrm{b}}}=b(\cos\phi_{\xi},\sin\phi_{\xi},0),\qquad{\boldsymbol{\mathrm{x}}}=\rho(\cos\phi,\sin\phi,0).$ (2)
@@
$A(\boldsymbol{\mathrm{x}})$ $=$ $E_{0}\frac{k}{iz_{0}z}\frac{1}{2\pi}\iint d^{2}\boldsymbol{\mathrm{b}}e^{ikS(\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})},$ (3)
where $k=2\pi/\lambda$ is the wavenumber and $S(\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})$ is the effective path length (eikonal) along the path from the source at $(0,-z_{0})$ to the observer’s position at $(\boldsymbol{\mathrm{x}},z)$ via a point $(\boldsymbol{\mathrm{b}},0)$ on the lens plane (see Fig. 2) Turyshev (2024)
$S(\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})$ $=$ $z_{0}+z+\frac{\boldsymbol{\mathrm{x}}^{2}}{2(z_{0}+z)}+\frac{z_{0}+z}{2z_{0}z}\Big{(}\boldsymbol{\mathrm{b}}-\frac{z_{0}}{z_{0}+z}\boldsymbol{\mathrm{x}}\Big{)}^{2}+\psi(\boldsymbol{\mathrm{b}})+{\cal O}(b^{4}),$ (4)
where $\psi(\boldsymbol{\mathrm{b}})$ is the gravitational phase shift that is acquired by the wave as it moves along its geodetic path from the source to the observer. For a monopole lens it has the form Turyshev and Toth (2021); Turyshev and Toth (2022):
@@
$A(\boldsymbol{\mathrm{x}})$ $=$ $\frac{E_{0}}{z_{0}+z}\exp\Big{[}ik\Big{(}z_{0}+z+\frac{\boldsymbol{\mathrm{x}}^{2}}{2(z_{0}+z)}\Big{)}\Big{]}\cdot F(\boldsymbol{\mathrm{x}}),$ (6)
where $F(\boldsymbol{\mathrm{x}})$ is the amplification factor that is given by the following diffraction integral Turyshev and Toth (2021); Turyshev (2024)
$F(\boldsymbol{\mathrm{x}})$ $=$ $\frac{ke^{ikr_{g}\ln 4k^{2}zz_{0}}}{i\tilde{z}}\frac{1}{2\pi}\iint d^{2}\boldsymbol{\mathrm{b}}\exp\Big{[}ik\Big{(}\frac{1}{2\tilde{z}}\Big{(}\boldsymbol{\mathrm{b}}-\frac{\tilde{z}}{z}\boldsymbol{\mathrm{x}}\Big{)}^{2}-2r_{g}\ln kb\Big{)}\Big{]},\qquad\tilde{z}=\frac{z_{0}z}{z_{0}+z}.~{}~{}~{}~{}$ (7)
@@
Benefiting from the axial-symmetry, we integrate over $d\phi_{\xi}$. Then, using the method of stationary phase Turyshev and Toth (2017), we integrate over $db$, and present (7) as below (see details in Turyshev (2024))
$F(\boldsymbol{\mathrm{x}})$ $=$ $\sqrt{2\pi kr_{g}}e^{i\varphi}J_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho\Big{)},$ (8)
where $\varphi=\sigma_{0}+kr_{g}\ln 2k(z_{0}+z)$ with $\sigma_{0}=-kr_{g}\ln(kr_{g}/e)-\frac{\pi}{4}$, see Turyshev and Toth (2017), and where $J_{0}$ is the Bessel function of the first kind Abramowitz and Stegun (1965).

### v3

@@ Abstract
We consider interstellar light transmission facilitated by gravitational lensing. We study axially-symmetric lensing configurations where the transmitter, lens, and receiver are nearly aligned. With the transmitter positioned in the focal region of the lens, we investigate the caustic formed by a diffraction-limited beam of light emitted by the transmitter. We evaluated the impact of the lens’s point spread function (PSF) on the received beam. We estimate the power delivered to a receiver at interstellar distances, evaluate the major noise sources, and determine the detection sensitivity in both the noise- and signal-dominated regimes. Considering realistic assumptions about the transmitter’s performance, we explore signal detection strategies enhanced by the spatial broadening of the received beam, a result of the PSF from the transmitting lens. We find that detection of the lensed optical signals from nearby stars may be done relying on established optical engineering technologies. Furthermore, a network of astronomical facilities capable of observations in multiple narrow spectral bands will benefit the search. Our results support the feasibility of interstellar power transmission via gravitational lensing, directly contributing to the ongoing optical SETI efforts.
@@ I Introduction
Interstellar power transmission is very challenging. Even for a collimated laser beam, the large distances involved result in a very small energy received. Consider, a diffraction-limited telescope with aperture $d_{\tt T}$ yielding the beam divergence of $\theta_{0}\simeq\beta(\lambda/d_{\tt T})=1.00\times 10^{-6}\,(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt T})$, where $\beta$ is a coefficient that depends on the type of light amplitude distribution and the definition of beam diameter, typically of $\beta\simeq{\cal O}(1)$. When the signal reaches the receiver at distance $z$, the beam is expanded to a spot with the radius of $\rho_{*}\simeq z(\lambda/d_{\tt T})\simeq 3.09\times 10^{11}\,{\rm m}~{}(z/10\,{\rm pc})(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt T})$. As a result, in the case of a free-space laser power transmission in the vacuum, a telescope with the aperture $d_{\tt R}$ receives only a small fraction of the transmit power $P_{\tt T}$ given as $P_{\tt R}=P_{\tt T}{\pi({\textstyle\frac{1}{2}}d_{\tt R})^{2}}/{\pi\rho_{*}^{2}}$ or
@@
which yields photon flux of only $Q_{\tt R}=(\lambda/hc)P_{\tt R}\simeq 1.32\times 10^{-5}$  phot/s for $\lambda=1\,\mu$m. These signals are weak and require both 1) precise and stable transmitter pointing toward us to within $\rho_{\tt L}/z\simeq{\rm few}\,\mu$rad and 2) us having instruments capable of detecting such signals against the challenging optical backgrounds that are present in our stellar neighborhood. Therefore, unless there is a significant effort to send us a message using a powerful, focused beam directly aimed at us, detecting faint and transient signals with optical SETI [^1] See details on the ongoing SETI efforts at https://en.wikipedia.org/wiki/Search_for_extraterrestrial_intelligence remains extremely challenging.
@@
This paper is organized as follows: In Section II, we present the wave-theoretical tools to describe light propagation in a gravity field. We discuss the caustic formed by the diffraction-limited beam of light. We consider diffraction of light in a spherically-symmetric gravitational field and consider the process of the interstellar power transmission. In Section III, we estimate the levels of the anticipated signals, introduce major noise sources, and evaluate signal-to-noise ratio (SNR). In Section IV, we discuss various detection strategies. Our conclusions are presented in Section V.
@@ II EM waves in a gravitational field
We consider a nearby star with mass $M_{\tt L}$, radius $R_{\tt L}$, and its Schwarzschild radius of $r_{g}=2GM_{\tt L}/c^{2}$. We assume a transmitter placed on the optical axis—the line connecting the solar system and this stellar gravitational lens, as illustrated in Fig. 1. We also assume that this transmitter is facing the lens and is positioned in the lens’ focal region at a distance $z_{0}>R^{2}_{\tt L}/2r_{g}=547.8\,(R_{\tt L}/R_{\odot})^{2}(M_{\odot}/M_{\tt L})$ AU from it, on the side opposite the solar system. Next, we assume that this is a CW laser optical transmitter that is capable of coherently illuminating an annulus around the lens with a radius of $b=\sqrt{2r_{g}z_{0}}>R_{\tt L}$, which is chosen to ensure the beam’s partial focusing at the receiver Turyshev (2024). To achieve this, a diffraction-limited transmitter must produce an annular beam of light. [^2] See Duocastella and Arnold (2012); Livneh et al. (2018) on the current efforts in producing annular laser beams. This beam should have the mean diverging angle of $b/z_{0}\simeq 7.16\times 10^{-6}\,({b}/{R_{\tt L}})(R_{\tt L}/{R_{\odot}})({650\,{\rm AU}}/{z_{0}})~{}{\rm rad}$ relative to the optical axis. As shown by Fig. 1, there is a finite-size caustic that is formed in the focal region of the lens. What determines the dimensions of this caustic? In the following discussion, we examine the geometry of the region where beam focusing occurs.
@@
Figure 1: Annular beam propagation in a thin lens approximation showing transmitter, lens, annulus, and receiver plane.
II.1 Lensing caustic
It is known that when a plane EM wave, originating at infinity, travels near a massive body, its wavefront experiences bending (see Figs. 2, 3 in Turyshev and Toth (2017)). In the weak field approximation of general relativity, this deflection angle is $\theta_{\tt gr}={2r_{g}}/{b}$, where $b$ is a light rays’s impact parameter. As a result, a massive celestial body acts as a lens by focusing the EM radiation (i.e., where the light rays with the same $b$ intersect the optical axis) at distance $z$, that is determined as
$\frac{b}{z}=\theta_{\tt gr}\qquad\Rightarrow\qquad z=\frac{b^{2}}{2r_{g}}\qquad{\rm or}\qquad b=\sqrt{2r_{g}z}.$ (2)
Clearly, the rays with different impact parameters will intersect the optical axis at different distances, which is consistent with lens’ spherical aberration. As a result, for an infinitely large incident plane wave coming from infinity, a monopole gravitational lens forms a caustic with a semi-infinite focal line that begins from $z_{0}=R_{\tt L}/2r_{g}$ Turyshev and Toth (2017).
In our case, the situation is reversed: the transmitter, placed in the lens’s focal region at a large but finite distance $z_{0}>R_{\text{L}}/2r_{g}$ from it, faces the lens and emits a finite-width beam of light towards it. What shape and dimensions will the caustic, forming in the focal region behind the lens, have?
To answer this question, we will use the geometrical optics, the small angles, and the thin lens approximations. When the light rays are coming from a source that is located at a large but finite distance from the lens, $z_{0}$, for them to intersect the lens plane with the same impact parameter $b$, they must be emitted at an angle $\beta=b/z_{0}$. As the rays travel through the thin lens plane, direction of their travel is affected by lens’ gravity. As a result, the angle at which the rays with the same $b$ will intersect the optical axis now is smaller and is given as $\alpha=\theta_{\tt gr}-b/z_{0}$, which has to be positive (see Turyshev (2024) for discussion of the impact of different values of $\alpha$ on focusing), yielding
$\frac{b}{z}=\frac{2r_{g}}{b}-\frac{b}{z_{0}}\qquad\Rightarrow\qquad b=\sqrt{2r_{g}\tilde{z}},\qquad{\rm where}\qquad\tilde{z}=\frac{z_{0}z}{z_{0}+z}.$ (3)
This expression implies that if the transmission distance, $z_{0}$, and the reception distance, $z$, are known, then the transmitter has to the lens plane illuminate at a particular impact parameter $b=\sqrt{2r_{g}\tilde{z}}$, as determined by (3).
With $\alpha<\theta_{\text{gr}}$, the distance $z=b/\alpha$, at which light rays with the same impact parameter $b$ intersect the optical axis (see Fig. 1), increases compared to the case of a plane wave originating from infinity, which was described by (2). In Turyshev and Toth (2019); Turyshev (2024), we have shown that the expression for $z$ given (2) is now modified to read
$\frac{b}{z}=\frac{2r_{g}}{b}\Big{(}1-\frac{b^{2}}{2r_{g}z_{0}}\Big{)}\equiv\alpha\qquad\Rightarrow\qquad z=\frac{b^{2}}{2r_{g}}\frac{1}{1-b^{2}/2r_{g}z_{0}}.$ (4)
We now consider a diffraction-limited beam of light leaving the telescope at a divergence angle of $\sim\lambda/d_{\tt T}$, where $d_{\tt T}$ is the aperture of a telescope used to transmit the signal. When the beam reaches the lens plane, it produces an annulus with the mean radius of $b$ and the thickness of $2\Delta b$ with $\Delta b=(\lambda/d_{\tt T})z_{0}\simeq 9.72\times 10^{7}(\lambda/1\,\mu{\rm m})(1\,{\rm m}/d_{\tt T})(z_{0}/650\,{\rm AU})\,{\rm m}\simeq 0.14R_{\odot}$. The entire transmitted energy is deposited into the annulus with the area of $A=4\pi b\Delta b=4\pi b(\lambda/d_{\tt T})z_{0}$. As the beam propagates further, it is affected by the lens’ spherical aberration.
By the time the beam reaches the focal region, located at an interstellar distance, the beam has expanded, forming a caustic, see Fig. 1. The caustic features an axially symmetric shape resembling a bicone, consisting of two conjoined cones that meet at their bases yet have slightly different heights. What are the diameter and the length of this caustic?
To investigate the dimensions of this caustic, we will use the thin lens approximation. Fig. 1 shows the beam with the mean impact parameter $b$. The divergence of this beam, due to optical diffraction, produces an annulus on the lens plane where the two most distant rays have impact parameters $b_{\pm}=b\pm\Delta b$.
Similar to (4), we can determine the angles $\alpha_{\pm}$ and the distances $z_{\pm}$ at which these rays intersect the optical axis:
$\frac{b_{\pm}}{z_{\pm}}=\frac{2r_{g}}{b_{\pm}}\Big{(}1-\frac{b^{2}_{\pm}}{2r_{g}z_{0}}\Big{)}\equiv\alpha_{\pm}\qquad\Rightarrow\qquad z_{\pm}=\frac{b^{2}_{\pm}}{2r_{g}}\frac{1}{1-b^{2}_{\pm}/2r_{g}z_{0}}.$ (5)
Taking into account that $\Delta b\ll b$, indeed
$\frac{\Delta b}{b}=\frac{(\lambda/d_{\tt T})z_{0}}{b}=0.14\Big{(}\frac{\lambda}{1\,\mu{\rm m}}\Big{)}\Big{(}\frac{1\,{\rm m}}{d_{\tt T}}\Big{)}\Big{(}\frac{z_{0}}{650\,{\rm AU}}\Big{)}\Big{(}\frac{R_{\tt L}}{b}\Big{)}\Big{(}\frac{R_{\odot}}{R_{\tt L}}\Big{)},$ (6)
we may develop all the relevant expressions to the first order in $\Delta b/b$.
First of all, we compute the width of the beam at the point $z$, where the central ray intersects the optical axis. As seen in Fig. 1, the diameter of this area is given by $d_{z}=2h=2\alpha_{+}(z_{+}-z)$. Using (4) and (5), we express this as
$d_{z}=2h=2b_{+}\Big{(}1-\frac{z}{z_{+}}\Big{)}=\frac{4\Delta b}{1-b^{2}/2r_{g}z_{0}}+{\cal O}\big{(}\Delta b^{2}\big{)}.$ (7)
Similarly, we determine the length of the caustic $\Delta z=z_{+}-z_{-}$ as it extends alone the optical axis
$\Delta z=z_{+}-z_{-}=\frac{4\Delta b}{1-b^{2}/2r_{g}z_{0}}\frac{z}{b}+{\cal O}\big{(}\Delta b^{2}\big{)}.$ (8)
With result for the impact parameter $b=\sqrt{2r_{g}\tilde{z}}$ from (3), we express $1-b^{2}/2r_{g}z_{0}=1-\tilde{z}/z_{0}=z_{0}/(z_{0}+z)$. Using this result in (9)–(10), and considering $\Delta b=(\lambda/d_{\text{t}})z_{0}$, we determine the diameter and length of the caustic as below
$d_{z}$ $=$ $\frac{4\lambda}{d_{\tt T}}(z_{0}+z)=8.25\Big{(}\frac{\lambda}{1\,\mu{\rm m}}\Big{)}\Big{(}\frac{1\,{\rm m}}{d_{\tt T}}\Big{)}\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}~{}{\rm AU},$ (9)
$\Delta z$ $=$ $\frac{zd_{z}}{\sqrt{2r_{g}\tilde{z}}}=\frac{d_{z}}{\theta_{\tt ER}}=16.29\Big{(}\frac{\lambda}{1\,\mu{\rm m}}\Big{)}\Big{(}\frac{1\,{\rm m}}{d_{\tt T}}\Big{)}\Big{(}\frac{z}{10\,{\rm pc}}\Big{)}^{2}\Big{(}\frac{M_{\odot}}{M_{\tt L}}\Big{)}^{\frac{1}{2}}\Big{(}\frac{650\,{\rm AU}}{z_{0}}\Big{)}^{\frac{1}{2}}~{}{\rm kpc},$ (10)
where $\theta_{\tt ER}$ is the radius of the Einstein ring as seen by the observer at $z$ (see (20) below for details).
Interestingly, the dimensions of the caustic created by a monopole lens for an annular beam are finite but very large. The diameter $d_{z}$ is a factor 2 larger than that for a free-space propagation. This is due beam broadening by the monopole gravitational lens. From a practical standpoint, increasing the transmitting telescope’s diameter would result in a more compact caustic, thereby enhancing the power density at the receiver thus aiding in signal detection.
Although light is present everywhere within the bicone-shaped caustic areas with dimensions of $(\Delta z,d_{z})$ given by (9)–(10), not all regions within it are of interest to us. In fact, as shown in Fig. 1, similarly to the scenario with a monopole lens and an incident plane wave Turyshev and Toth (2017), the propagation of an annular diffraction-limited beam gives rise to four distinct regions: shadow, geometrical optics, and both weak and strong interference regions, namely: i) Since there is no light with impact parameters that fall outside the range estublished by the diffraction limit, $b\in[b_{-},b_{+}]$, the corresponding areas form a shadow region behind the lens, where no light from the transmitter is present; ii) The regions through which only one ray from the source passes at each point (i.e., outside the interference region set by the biconical shape in Fig. 1) correspond to the regions of geometric optics; iii) As we enter the bicone-shaped area, we move into the zone of weak interference, where at least two rays from opposite sides of the annulus intersect. Here, the light intensity begins to increase, marking it as the region of weak interference; iv) As we approach closer to the optical axis, in the immediate vicinity of it, at separations $r_{\tt GL}\leq r_{g}(z/z_{0})\simeq 9.37\times 10^{6}(M_{\tt L}/M_{\odot})(z/10\,{\rm pc})(650\,{\rm AU}/z_{0})$ m, we enter the region of strong interference. Here, near the optical axis, an intensely concentrated beam is present. Proper description of the EM field in the interference regions requires a wave-theoretical treatment, which we discuss next.
II.2 Light amplification
@@
At a high level, we know that when electromagnetic (EM) waves travel in the vicinity of a stellar gravitational lens, its gravity causes the waves to scatter and diffract Turyshev (2017); Turyshev and Toth (2017). Due to the spherical aberration inherent in a monopole lens, the beam will not focus at a single point. Instead, the lens’s gravitational field will cause the light to converge within the finite-size caustic forming the focal region in the vicinity of the optical axis, as discussed in Turyshev (2024) and shown in Fig. 1. The properties of this caustic are driven by the transmission annulus, with the mean radius of $b$ and the thickness of $2\Delta b\sim 2(\lambda/d_{\tt T})z_{0}$. According to (3), for a properly chosen transmitter position, $z_{0}$, the transmitted light will be deposited in the solar system at distance $z\gg z_{0}$, where the beam will be compressed and partially focused. A solar system observer, looking back at the lensing star, will see the transmitted signal in the form of the Einstein ring around the lens. Compared to a free-space propagation (1), such gravitational focusing results in a major increase of the power density of the EM field deposited in the receiver plane Turyshev (2024). Here, we consider the physics of this process.
@@
We introduce a lens-centric cylindrical coordinate system $(\rho,\phi,z)$ with its $z$-axis oriented along the unperturbed direction of the incident wave’s propagation (see Fig. 2), given by a unit vector $\boldsymbol{\mathrm{k}}$. We also introduce a light ray’s impact parameter, $\boldsymbol{\mathrm{b}}$, with $\boldsymbol{\mathrm{b}}\perp\boldsymbol{\mathrm{k}}$. In this coordinate system, a ray of light is emitted from a point with coordinates $(\boldsymbol{\mathrm{x}}^{\prime},-z_{0})$, while it is received in the receiver plane at a point with coordinates $(\boldsymbol{\mathrm{x}},z)$ and
${\boldsymbol{\mathrm{x}}}^{\prime}=\rho^{\prime}(\cos\phi^{\prime},\sin\phi^{\prime},0),\qquad{\boldsymbol{\mathrm{b}}}=b(\cos\phi_{\xi},\sin\phi_{\xi},0),\qquad{\boldsymbol{\mathrm{x}}}=\rho(\cos\phi,\sin\phi,0).$ (11)
@@
$A(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})$ $=$ $E_{0}\frac{k}{iz_{0}z}\frac{1}{2\pi}\iint d^{2}\boldsymbol{\mathrm{b}}e^{ikS(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})},$ (12)
where $k=2\pi/\lambda$ is the wavenumber and $S(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})$ is the effective path length (eikonal) along the path from the source at $(\boldsymbol{\mathrm{x}}^{\prime},-z_{0})$ to the observer’s position at $(\boldsymbol{\mathrm{x}},z)$ via a point $(\boldsymbol{\mathrm{b}},0)$ on the lens plane (see Fig. 2 and Turyshev (2024))
$S(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{b}},\boldsymbol{\mathrm{x}})$ $=$ $\sqrt{(\boldsymbol{\mathrm{b}}-\boldsymbol{\mathrm{x}}^{\prime})^{2}+z_{0}^{2}}+\sqrt{(\boldsymbol{\mathrm{b}}-\boldsymbol{\mathrm{x}})^{2}+z^{2}}+\psi(\boldsymbol{\mathrm{b}})=$ (13)
$=$ $z_{0}+z+\frac{(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{\prime})^{2}}{2(z_{0}+z)}+\frac{z_{0}+z}{2z_{0}z}\Big{(}\boldsymbol{\mathrm{b}}-\frac{z_{0}}{z_{0}+z}\Big{(}\boldsymbol{\mathrm{x}}+\frac{z}{z_{0}}\boldsymbol{\mathrm{x}}^{\prime}\Big{)}\Big{)}^{2}+\psi(\boldsymbol{\mathrm{b}})+{\cal O}(b^{4}),$
where $\psi(\boldsymbol{\mathrm{b}})$ is the gravitational phase shift that is acquired by the wave as it moves along its geodetic path from the source to the observer. For a monopole lens it has the form Turyshev and Toth (2021a); Turyshev and Toth (2022):
@@
$A(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})$ $=$ $\frac{E_{0}}{z_{0}+z}\exp\Big{[}ik\Big{(}z_{0}+z+\frac{(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{\prime})^{2}}{2(z_{0}+z)}\Big{)}\Big{]}\cdot F(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}}),$ (15)
where $F(\boldsymbol{\mathrm{x}})$ is the amplification factor that is given by the following diffraction integral Turyshev and Toth (2021a); Turyshev (2024)
$F(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})$ $=$ $\frac{ke^{ikr_{g}\ln 4k^{2}zz_{0}}}{i\tilde{z}}\frac{1}{2\pi}\iint d^{2}\boldsymbol{\mathrm{b}}\exp\Big{[}ik\Big{(}\frac{1}{2\tilde{z}}\Big{(}\boldsymbol{\mathrm{b}}-\frac{\tilde{z}}{z}\Big{(}\boldsymbol{\mathrm{x}}+\frac{z}{z_{0}}\boldsymbol{\mathrm{x}}^{\prime}\Big{)}\Big{)}^{2}-2r_{g}\ln kb\Big{)}\Big{]},\qquad\tilde{z}=\frac{z_{0}z}{z_{0}+z}.~{}~{}~{}~{}$ (16)
@@
Benefiting from the axial-symmetry, we integrate over $\phi_{\xi}$ from 0 to $2\pi$. Then, to integrate over the impact parameter, $b$, we use the method of stationary phase Turyshev and Toth (2017) and present (16) as below (see details in Turyshev (2024))
$F(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})$ $=$ $\sqrt{2\pi kr_{g}}e^{i\varphi}J_{0}\Big{(}k\frac{\sqrt{2r_{g}\tilde{z}}}{z}\rho(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})\Big{)},\qquad~{}~{}\rho(\boldsymbol{\mathrm{x}}^{\prime},\boldsymbol{\mathrm{x}})=\big{|}\boldsymbol{\mathrm{x}}+\frac{z}{z_{0}}\boldsymbol{\mathrm{x}}^{\prime}\big{|},$ (17)
where the impact parameter for which the phase is stationery was determined to be $b=\sqrt{2r_{g}\tilde{z}}$. Also, $\varphi=\sigma_{0}+kr_{g}\ln 2k(z_{0}+z)$ with $\sigma_{0}=-kr_{g}\ln(kr_{g}/e)-\frac{\pi}{4}$, see Turyshev and Toth (2017), and where $J_{0}$ is the Bessel function of the first kind Abramowitz and Stegun (1965).